# MIMIR: Masked Image Modeling for Mutual Information-based Adversarial Robustness

### v2

@@ Abstract
Vision Transformers (ViTs) achieve superior performance on various tasks compared to convolutional neural networks (CNNs), but ViTs are also vulnerable to adversarial attacks. Adversarial training is one of the most successful methods to build robust CNN models. Thus, recent works explored new methodologies for adversarial training of ViTs based on the differences between ViTs and CNNs, such as better training strategies, preventing attention from focusing on a single block, or discarding low-attention embeddings. However, these methods still follow the design of traditional supervised adversarial training, limiting the potential of adversarial training on ViTs. This paper proposes a novel defense method, MIMIR, which aims to build a different adversarial training methodology by utilizing Masked Image Modeling at pre-training. We create an autoencoder that accepts adversarial examples as input but takes the clean examples as the modeling target. Then, we create a mutual information (MI) penalty following the idea of the Information Bottleneck. Among the two information source inputs and corresponding adversarial perturbation, the perturbation information is eliminated due to the constraint of the modeling target. Next, we provide a theoretical analysis of MIMIR using the bounds of the MI penalty. We also design two adaptive attacks when the adversary is aware of the MIMIR defense and show that MIMIR still performs well. The experimental results show that MIMIR improves (natural and adversarial) accuracy on average by 4.19% on CIFAR-10 and 4.64% on ImageNet-1K, compared to baselines. On Tiny-ImageNet, we obtained improved natural accuracy of 2.99% on average and comparable adversarial accuracy. Our code and trained models are publicly available [^1] https://github.com/xiaoyunxxy/MIMIR.
Index Terms:
Mask Image Modeling, Adversarial Robustness, Vision Transformer, Mutual Information
@@ 1 Introduction
While Vision Transformers [28] and the variants [51, 24] have achieved great advances, these attention-based technologies fail in being more robust than CNNs against evasion attacks in common threat models [6, 1, 57, 87, 35]. Evasion attacks [11, 71] (also known as adversarial attacks) refer to fooling well-trained deep models by adding human-imperceptible perturbations to inputs. To build robust transformer models, adversarial training is considered one of the most successful methods [87, 57, 25], in which the general idea is to use adversarial examples for training. However, adversarial training has limitations due to its expensive computational cost to generate adversarial examples. Significant efforts have been made to decrease the adversarial training cost [67, 86, 80, 69], but the nature of attention-based transformer introduces new challenges for accelerating adversarial training. For example, transformers lack inductive biases [28], including locality, two-dimensional neighborhood structure, and translation equivariance. These biases are inherent to CNNs as prior knowledge but not applicable to ViTs [28]. The lack of such prior knowledge makes ViTs require a large amount of training data to generalize well, which means generating more adversarial perturbations for adversarial training.
Due to the differences between CNN and ViTs, previous works try to design new methodologies for adversarial training when considering the robustness of ViTs. Gu et al. [35] proposed Smoothed Attention to improve the robustness against patch-wise attacks by blocking attention to individual patches. Attention-Guided Adversarial Training (AGAT) [87] can efficiently drop image embeddings with lower attention and keep the higher ones after each layer of self-attention, accelerating adversarial training. Mo et al. [57] found that the robustness of ViTs can be improved by randomly masking perturbations or masking gradients from some attention blocks. In addition to new methodologies, some work has also tried to find better training strategies for the robustness of ViTs. The authors of [57] observed beneficial behaviors for adversarial training of ViTs, including fine-tuning on clean pre-trained models, using gradient clipping, and combining SGD optimizer and piecewise learning rate scheduler. Bai et al. [6] discussed that adversarial examples can be too difficult for ViTs to learn when combined with strong data augmentation, such as Randaugment [21], CutMix [95], and MixUp [97]. Then, they proposed a warmup strategy for augmentations to alleviate this issue. They progressively increase the distortion magnitudes of Randaugment and the probability of applying MixUp or CutMix. Debenedetti et al. [25] also found that strong data augmentations harm adversarial training. They proposed to use basic data augmentations (random-resize-and-crop + horizontal-flipping + color-jitter) and high weight decay for better performance.
However, previous research focuses on adversarial training at the fine-tuning stage or lacks rigorous explanations, limiting the potential for improving adversarial training. For example, AGAT [87] accelerates adversarial training by dropping a constant ratio of low-attention image embeddings after every layer. Overall, they reduce 40% computation of the entire model. Designing a method that discards more image embeddings is challenging because the accuracy of clean inputs also decreases when a larger number of embeddings are dropped. We investigate a novel approach where we conduct adversarial training at the pre-training stage. Our motivation is to discard a much larger ratio of image embeddings at the pre-training stage, further accelerating adversarial training. In addition, we show a very intuitive analysis of the training process (i.e., decreasing MI between adversarial inputs and latent representation) with theoretical proof, which ensures a good performance. The above motivations yield a self-supervised learning task suitable for Masked Image Modeling (MIM). MIM refers to learning from unlabeled data through masked inputs to recover the original information, such as recovering the original visual tokens [7] or original image pixels [36] based on the masked image patches, masking out a portion of the input sequence and then predicting the feature of the masked regions [85]. Usually, MIM methods can mask out a much larger ratio of inputs by self-supervised pre-training. For example, MAE [36] masks out 75% of image patches at the beginning of the forward pass. The reason is that MIM intends to learn more discriminative information, so masking out the discriminative part of images and recovering it gives a more difficult reconstructing task [81]. The reconstructing task is more difficult when applying an adversarial perturbation to the original image, as the perturbation corrupts the natural features in clean images. In this way, the model learns more discriminative information, which is also robust to adversarial perturbation. Consequently, adversarial training can be both efficient and effective.
Specifically, we propose Masked Image Modeling for Mutual Information-based Adversarial Robustness (MIMIR) as a pre-training method, where we use adversarial examples as inputs and clean examples as reconstructing objectives. As illustrated in Figure 1, MIMIR [^2] MIMIR is a figure in Norse mythology renowned for his knowledge and wisdom. consists of an encoder and a decoder following the design of MAE [36]. The encoder is a ViT, which has the same architecture as the model to be fine-tuned but without the final classification layer. The decoder consists of a few lightweight ViT blocks that reconstruct the original clean image from the latent features generated by the encoder. At pre-training, adversarial examples are separated into image patches. A large ratio of image patches is masked, so the encoder only processes a small part of adversarial patches. At fine-tuning, the decoder is discarded. The encoder is combined with the initialized classification layer and applied to complete images for classification tasks. As we use latent representation from adversarial examples to reconstruct clean images, this represents a more difficult reconstruction task. In addition, we theoretically analyze the behaviors of our encoder and decoder following the idea of Information Bottleneck [72]. The MI between adversarial inputs and latent representation will decrease concerning the accuracy of adversarial examples while pre-training. This motivates us to directly add an MI penalty (between adversarial inputs and latent representation) into the loss function of MIMIR at pre-training. This new pre-training loss further improves performance at the fine-tuning stage (see Table VII). We perform comprehensive experiments on CIFAR-10 [46], Tiny-ImageNet [47], and ImageNet-1K [26] and compare to state-of-the-art results. The results demonstrate that MIMIR significantly improves the performance on both natural and adversarial accuracy. Finally, previous works [25, 57, 6] claim that strong data augmentations (Randaugment [21], CutMix [95], and MixUp [97]) are unsuccessful in adversarial training. However, we show in Table VI that strong augmentations actually improve the natural and adversarial accuracy with the longer training schedule. Our key contributions are summarized as follows:
  • We propose a novel and efficient defense against adversarial attacks on ViTs - MIMIR, which discards 75% of image patches and then reconstructs the complete clean image under a self-supervised framework.
  • We provide a theoretical analysis on using adversarial examples and MI penalty in MIMIR. We show that MI between adversarial examples and learned latent representation should be decreased.
@@
  • We systematically evaluate MIMIR using three datasets and three transformer architectures with multiple model scales under various adversarial attacks.
@@
  • We show MIMIR is resistant to two adaptive attacks when the adversary is aware of MIMIR. First, we build a PGD-feature attack inspired by [65], which directly increases the distance between features extracted from clean and adversarial examples. Then, we build a PGD MI attack that uses cross entropy and the MI penalty as learning objectives to find adversarial examples.
Figure 1: The structure of MIMIR. First, we generate adversarial perturbation $\delta$ and apply it to the input image. Then, following the original ViT, we use separated image patches as input for the autoencoder. We use the decoder output, i.e., $x_{re}$, and the clean input image, i.e., $x$, to calculate the loss. Finally, we use MI ($I(x+\delta,z)$) as an additional penalty in the pre-training loss. After pre-training, the encoder plus a classification layer is fine-tuned for classification tasks.
2 Background
2.1 Adversarial Attack
@@
  1. 1.
    The generated perturbation $\delta$ can successfully mislead the network by maximizing (e.g., PGD attack [55]):
@@
[…]
  2. 2.
    The generated adversarial examples are as similar as possible to the original clean examples by limiting $\delta$ to a relatively small domain:
@@
When using the above attacks to generate adversarial examples for training, the learning objective is:
@@
2.2 Masked Image Modeling - MIM
MIM refers to a self-supervised pre-training framework that aims to reconstruct pre-defined targets, such as discrete tokens [7], raw RGB pixels [36, 49], or features [85]. The final goal is to use the pre-trained model as a starting point for downstream fine-tuning. The downstream tasks include, for instance, classification and object detection. More specifically, to build a high-performance ViT $f_{e}$ without a classification layer, we consider $f_{e}$ as an encoder to extract discriminative input features. Then, we design a lightweight decoder $f_{d}$, which uses the output of $f_{e}$ as its input. The goal of $f_{d}$ is to reconstruct the original inputs (let us consider MAE [36] as an example). The aim is to decrease the distance between the input $x$ and $x_{re}=f_{d}\circ f_{e}(x)$. After the encoder $f_{e}$ and decoder $f_{d}$ are trained, we use $f_{e}$ plus a manually initialized classification layer as the starting point of fine-tuning.
2.3 Mutual Information - MI
MI measures the mutual dependence between two random variables, $X$ and $Y$. It quantifies the amount of information contained in one random variable about another random variable or the reduced uncertainty of a random variable when another random variable is known. It can be written as:
$I(X,Y)=\int_{\mathcal{Y}}\int_{\mathcal{X}}P_{(X,Y)}(x,y)\log\left(\frac{P_{(X,Y)}(x,y)}{P_{(X)}(x)P_{(Y)}(y)}\right),$ (4)
where $P_{(X,Y)}$ is the joint probability density function of $X$ and $Y$. $P_{(X)}$ and $P_{(Y)}$ are the marginal probability density function of $X$ and $Y$. In addition, MI can be equivalently expressed as:
$I(X,Y)$ $=H(X)-H(X|Y).$ (5)
However, quantifying MI is difficult in practice, as it is difficult to precisely estimate $P_{X,Y}$ or $P_{X}$ and $P_{Y}$ in high dimensional space [31]. DIB [94] suggested using the matrix-based Renyi’s $\alpha$-order entropy $I_{\alpha}$ [66, 93] to estimate MI, which avoids density estimation and variational approximation. An alternative way is the Hilbert-Schmidt Independence Criterion (HSIC) [34], which is a kernel-based dependence measure defined in a reproducing kernel Hilbert space (RKHS) and usually used as a surrogate of MI. Details about definitions and empirical estimators of $I_{\alpha}$ and HSIC are provided in Appendix G. In this paper, we use both $I_{\alpha}$ and HSIC as our MI measurements.
2.4 Information Bottleneck - IB
@@
The IB concepts were first proposed in [72] and further developed for deep learning in [73, 68]. It has been widely used to improve adversarial robustness [90, 2, 94, 84]. IB describes the generalization of a deep network in two phases: 1) empirical error minimization (ERM) and 2) representation compression [68]. For a network with input $x$ and label $y$, there is an intermediate representation $t_{l}$ for each layer $l$, i.e., the output of $l$-th layer. The IB principle aims to keep more relevant information in $t_{l}$ about target $y$ while decreasing the irrelevant information about input $x$. The information between intermediate representation $t_{l}$ and input $x$ or label $y$ is quantified by MI, denoted by $I(\cdot)$. During neural network training, in the ERM phase, the model increases shared information between $t_{l}$ with respect to both $x$ and $y$. Afterward, in the compression phase, the model decreases information contained in $t_{l}$ about $x$ but preserves (or even increases) information about $y$. The reduction of $I(x,t_{l})$ can be interpreted as a way of reducing noise or compressing irrelevant or redundant features in $x$. At the end of the training, the model strikes a trade-off that maximizes $I(y,t_{l})$ and minimizes $I(x,t_{l})$. Formally, the IB minimizes the following Lagrangian:
$\mathcal{L}=I(x,t_{l})-\beta I(y,t_{l}),$ (6)
where $\beta$ is a Lagrange multiplier that controls the trade-off between predicting $y$ and compressing $x$.
@@ 3.1 Threat Model
We assume the adversary intends to create adversarial examples from clean images. The adversary has the strongest ability, i.e., the white box, to access the target model. Then, adversarial examples can be created according to model architectures, parameters, the gradients of loss function, and datasets. In addition, the adversary is also aware of potential defenses. An adversary can design new attacks for a specific model according to the details of our defense method.
From the defender’s perspective, the main goal is to train a robust model against potential adversarial attacks. The defender considers the following four objectives:
  • The performance on clean data should not decrease significantly. We allow a small drop of clean accuracy for a trade-off in robustness.
  • The accuracy of models trained by MIMIR should be significantly higher than for models without defenses under various adversarial attacks.
  • Deploying MIMIR should be efficient and scalable to large datasets such as ImageNet-1K [26].
  • MIMIR should be robust against adaptive attacks.
@@ 3.2 Design Intuition
MIM has been proved useful as a pre-training method for ViTs on various tasks [79, 36, 85, 7]. To train a powerful model by MIM, one can mask out a part of the foreground of inputs and reconstruct it using the model. Masking out the foreground instead of the background reduces discriminative information in visible information to the model. Reconstructing foreground parts is harder than background and helps the model learn more discriminative information [81]. Inspired by this phenomenon, we aim to build a more difficult task by adding adversarial perturbation to clean inputs. The adversarial perturbation increases the distance between the input and reconstruction target. If we can reconstruct clean inputs from adversarial examples, it means the features learned by the model are robust against adversarial attacks. In other words, we want the encoder to learn a latent representation that does not carry information concerning adversarial inputs while enabling the decoder to reconstruct the clean image. Note that simply masking out all foreground does not make sense, as it would not be possible to reconstruct meaningful content. From the perspective of IB, there is a bottleneck between the encoder and decoder. Our adversarial pre-training task contains two information sources: the clean data $x$ and the adversarial perturbation $\delta$. As the information flows through the bottleneck, adversarial information from $\delta$ is eliminated. The natural information from $x$ is maintained because of the constraint when reconstructing the target $x$.
3.3 Design Details
Autoencoder. MIMIR follows the general design of MAE [36], which consists of an encoder $f_{e}$ and a decoder $f_{d}$. Like all other autoencoders, the encoder extracts discriminative features $z$ from inputs $x$. The decoder reconstructs original inputs according to the discriminative features. Following ViT [28], the input $x$ is separated into non-overlapping image patches. Then, we randomly mask out a part of the patches and use the remaining patches as inputs for the following process in the encoder. This random masking process uses uniform distribution to prevent potential sampling bias, such as all foreground information being masked, as it is not possible to find the reconstruction target then. Therefore, we hope to keep a part of the foreground as a hint for reconstruction. The information of masked content is recorded as mask tokens $m$, which are not used by the encoder but will be used later by the decoder. Each token is a learned vector that indicates the presence of a masked patch to be predicted. The mask token is shared by all inputs of the decoder. Like unmasked patches, mask tokens are also assigned corresponding positional embeddings to be in the correct location in the reconstructed image. We emphasize that mask tokens are not used in the encoder part.
To train a ViT, we use the same transformer blocks as ViT to build the encoder. The encoder only processes the visible patches, making training much more efficient. When converting to other architectures, such as ConVit [24], we use corresponding transformer blocks to build the encoder. The decoder accepts the encoded visible image patches and mask tokens as inputs. The decoder is built using the same transformer blocks as the encoder instead of using ViT [28] transformer blocks for all. Then, the decoder is followed by a fully connected layer, which outputs the same number of patches as the original image.
Adversarial pre-training target. The training target is to extract discriminative features from visible image patches by the encoder and then reconstruct the invisible patches by the decoder. Therefore, we need a differentiable measurement to quantify the distance between the original image and the reconstructed results. Following the original MAE [36], this distance is measured by Mean Squared Error (MSE). To create a more difficult reconstruction task, we apply adversarial perturbation $\delta$ on the inputs of the encoder. Thus, the adversarial perturbation is also masked along with the image when inputted into the encoder. The decoder reconstructs the original clean inputs by using the latent features $z$ extracted from adversarial examples. The outputs of decoder $x_{re}$ and $x$ are used to calculate the MSE loss, which is further used to optimize the model. Note that the reconstruction differs from the original MAE; we do not use the encoder inputs as reconstruction targets. Formally, the pre-training process (described in Algorithm 1) can be written as follows:
$z$ $=f_{e}(x+\delta),$ (7)
$x_{re}=f_{d}(z),$
$loss_{\text{mse}}$ $=\mathcal{L}_{\text{mse}}(x,x_{re}).$
MI as penalty. Inspired by IB, we show in Section 3.4 that MI between latent representation and adversarial examples decreases with respect to the accuracy on adversarial example, i.e., $I(x+\delta,z)$ is decreasing while training. Motivated by this finding, we directly use $I(x+\delta,z)$ as a penalty in our final loss function:
where $\lambda$ is a regularizer for the MI penalty. We use $I(x+\delta,z)$ instead of $I(x,z)$ as a penalty. This is because $x\to x+\delta\to z$ follows the Markov chain since $z$ is extracted from $x+\delta$. According to Data Processing Inequality (DPI) [8], $I(x,z)\leq I(x+\delta,z)$. Minimizing $I(x,z)$ in loss equals minimizing a lower bound of $I(x+\delta,z)$.
Algorithm 1 Masked Adversarial Pre-Training
@@ 3.3 Design Details
Generating adversarial examples. To conduct the adversarial pre-training, we need an attack that finds proper adversarial perturbation $\delta$. As the autoencoder does not provide classification outputs, it is not possible to directly use existing adversarial attacks, such as PGD [55]. Still, we can easily design a new algorithm to find $\delta$ by maximizing $loss_{\text{mse}}$ in Eq. (7). As the feature $z$ is extracted from only visible image patches, we only attack the visible patches. We do not add any perturbation to mask tokens since the outputs of the autoencoder are only impacted by visible patches. Then, the adversarial pre-training learning objective can be written as:

### v3

@@ Abstract
Vision Transformers (ViTs) achieve excellent performance in various tasks, but they are also vulnerable to adversarial attacks. Building robust ViTs is highly dependent on dedicated Adversarial Training (AT) strategies. However, current ViTs’ adversarial training only employs well-established training approaches from convolutional neural network (CNN) training, where pre-training provides the basis for AT fine-tuning with the additional help of tailored data augmentations. In this paper, we take a closer look at the adversarial robustness of ViTs by providing a novel theoretical Mutual Information (MI) analysis in its autoencoder-based self-supervised pre-training. Specifically, we show that MI between the adversarial example and its latent representation in ViT-based autoencoders should be constrained by utilizing the MI bounds. Based on this finding, we propose a masked autoencoder-based pre-training method, MIMIR, that employs an MI penalty to facilitate the adversarial training of ViTs. Extensive experiments show that MIMIR outperforms state-of-the-art adversarially trained ViTs on benchmark datasets with higher natural and robust accuracy, indicating that ViTs can substantially benefit from exploiting MI. In addition, we consider two adaptive attacks by assuming that the adversary is aware of the MIMIR design, which further verifies the provided robustness. Our code and trained models are publicly available at the anonymous link. [^1] https://anonymous.4open.science/r/MIMIR-5444
@@ 1 Introduction
Adversarial training (Madry et al. 2018) is one of the standard techniques to build robust models. Regarding adversarial training for ViTs, early works focus on investigating the unique attention mechanism of ViTs. There, robustness can be improved by randomly dropping gradient according to attention (Mo et al. 2022) or improving training efficiency by dropping low-attention image embeddings (Wu et al. 2022). The majority of recent works focus on improving the ViTs’ adversarial training based on experience with CNNs, given AT’s success in building robust CNNs. Looking at the adversarial training recipe differences between ViTs and CNNs, strong data augmentation, e.g., Randaugment (Cubuk et al. 2020), CutMix (Yun et al. 2019), and MixUp (Zhang et al. 2018), show positive influence in CNN but are ineffective for ViTs AT and sometimes even reduce the AT performance of ViTs (Bai et al. 2021; Singh, Croce, and Hein 2023). To mitigate this, the distortion magnitudes of Randaugment and the probability of applying MixUp or CutMix need to be progressively increased (Bai et al. 2021). Similar problems of strong data augmentations are also observed by (Debenedetti, Sehwag, and Mittal 2023), where basic data augmentations, i.e., random-resize-and-crop, horizontal-flipping, and color-jitter with high weight decay are shown to be promising to improve ViT adversarial training.
Another common practice in ViT adversarial training is to leverage pre-trained models, which has also been extensively explored (Mo et al. 2022; Singh, Croce, and Hein 2023). Fine-tuning on natural pre-trained models could benefit the adversarial robustness of ViTs. For instance, recently, AdvXL (Wang et al. 2024) introduced an efficient and effective adversarial training strategy for giant models on web-scale datasets. However, as the current ViT adversarial training strategies are mainly inherited from CNN AT, the key factor that drives the success of pre-training has not been completely revealed. For instance, pre-training with ImageNet-21K and SimMIM (Xie et al. 2022) performs similarly to the model trained from scratch, and even the CLIP (Radford et al. 2021) pre-trained model shows a performance drop (Liu et al. 2024).
In this work, we aim to provide an in-depth analysis of ViT’s robustness using MI, focusing on self-supervised pre-training. Information Bottleneck (IB) has been previously used in supervised learning to analyze adversarial robustness, where a joint objective of simultaneously minimizing the MI between inputs and latent features while maximizing the MI between labels and latent features is introduced to mitigate the impact of the adversarial noise in the inputs. Regarding the task of ViT adversarial training, we provide a novel theoretical analysis for the self-supervised autoencoder, showing that the self-supervised ViT benefits from decreasing the MI between inputs and latent features. Based on this finding, we propose a theoretically grounded adversarial pre-training method, Masked Image Modeling for Mutual Information-based Adversarial Robustness (MIMIR). Figure 1 provides an illustrative diagram of MIMIR.
MIMIR substantially improves the adversarial robustness of ViT models as demonstrated by comprehensive experiments on ImageNet-1K (Deng et al. 2009), Tiny-ImageNet (Le and Yang 2015), and CIFAR-10 (Krizhevsky, Hinton et al. 2009). We also provide an experimental comparison with the state-of-the-art, showing the effectiveness of MIMIR in improving both natural and adversarial prediction accuracy (see Section 4.2).
Our main contributions are:
  • By revising the current ViT AT strategies, we point out that ViT adversarial pre-training methods compromise natural accuracy and lack systematic analysis. To this end, we provide a theoretical analysis using adversarial examples and the Mutual Information penalty. The MI bounds motivate us to decrease MI between adversarial examples and learned latent representation.
@@
  • Based on the findings of our analysis, we propose an efficient self-supervised defense MIMIR against adversarial attacks on ViTs. We systematically evaluate MIMIR using three datasets and multiple model scales under various adversarial attacks, demonstrating its superiority.
@@
  • We show MIMIR is resistant to two adaptive attacks, where the adversary is aware of MIMIR design. We build a PGD-fea attack that increases the distance between features extracted from natural and adversarial examples. Moreover, we build a PGD-MI attack that includes the MI penalty as the learning objective to find adversarial examples.
Figure 1: Diagram illustrating the working mechanism of MIMIR. In the pre-training, adversarial perturbations $\delta$ are generated and added to natural images $x$ to create adversarial images $x+\delta$, as shown on the left. Patches of generated adversarial images are fed as training data to the ViT-based autoencoder, where the output of decoder $x_{re}$ and the natural input image $x$ are used to calculate the loss. In particular, we propose using MI ($I(x+\delta,z)$) as an additional penalty to compose the pre-training loss of MIMIR, as shown in the bottom right. After pre-training, a trained encoder is combined with a randomly initialized classification layer as the final model that is further fine-tuned for classification tasks.
2 Background and Related Work
@@
The generated adversarial examples are as similar as possible to the original natural examples by limiting $\delta$ to a relatively small domain:
@@
When using the above adversarial examples to conduct adversarial training, the learning objective is:
@@
Information Bottleneck (IB). The IB concepts were first proposed in (Tishby, Pereira, and Bialek 2000) and further developed for deep learning in (Tishby and Zaslavsky 2015; Shwartz-Ziv and Tishby 2017). It has been widely used to improve adversarial robustness (Xu, Perin, and Picek 2023; Alemi et al. 2017; Yu, Yu, and Príncipe 2021; Wang et al. 2021). IB describes the generalization of a deep neural network in two phases: 1) empirical error minimization (ERM) and 2) representation compression. For a network with input $x$ and label $y$, there is an intermediate representation $t_{l}$ for each layer $l$, i.e., the output of $l$-th layer. The IB principle aims to keep more relevant information in $t_{l}$ about target $y$ while decreasing the irrelevant information about input $x$. The information between intermediate representation $t_{l}$ and input $x$ or label $y$ is quantified by MI, denoted by $I$.
Adversarial Robustness of ViTs. PGD (Madry et al. 2018) adversarial training is considered one of the most effective defenses that can withstand adaptively designed attacks (Athalye, Carlini, and Wagner 2018). Due to the difference between CNNs and ViTs, there have been substantial recent efforts to explore new adversarial training approaches for ViTs. EasyRobust (Mao et al. 2022) evaluates the robustness of ViTs using PGD adversarial training. (Mo et al. 2022; Bai et al. 2021; Debenedetti, Sehwag, and Mittal 2023) demonstrate the classical training recipe for CNN is not suitable for ViTs and present new adversarial training strategies for ViTs by evaluating different hyperparameters and data augmentation policies. (Mo et al. 2022) also finds that pre-training with natural data provides better robustness after adversarial fine-tuning. (Singh, Croce, and Hein 2023) explores the effect of strong pre-training and replacing the patch embedding in ViTs with convolutional blocks. AdvXL (Wang et al. 2024) shows an outstanding performance of giant models and web-scale data with adversarial pre-training and fine-tuning. However, (Liu et al. 2024) provides a different view that pre-training with ImageNet-21K and SimMIM (Xie et al. 2022) performs similarly to the model trained from scratch, and the CLIP pre-trained model has a performance drop. Unlike existing works, we provide a formal and in-depth analysis leveraging MI, based on which we introduce a new adversarial pre-training scheme.
@@ 3.1 Threat Model
Adversary’s Goal. The attacker aims to fool the trained model by decreasing the overall classification accuracy (non-targeted) or forcing the model to recognize an input as a specific target (targeted). The image perturbations are limited by the $l_{\infty}$ norm for invisibility.
Adversary Knowledge. The attacker has white-box access to the model, including training data, architectures, hyperparameters, and model weights. The attacker can implement iterative attacks and unlimited queries to update adversarial examples multiple times in white and black-box settings. Adversarial examples can be created according to model architectures, parameters, the gradients of loss function, and datasets. The adversary can adaptively design new attacks for a specific model according to the design details of the defense method.
Defender’s Goal. The defender’s main goal is to train a robust model against potential adversarial attacks and prevent the performance of natural data from decreasing significantly. In addition, the defense method should be robust against adaptive attacks specifically designed for it.
@@ 3.2 Design Intuition
MIM has been proved useful as a pre-training method for ViTs on various tasks (Vincent et al. 2010; He et al. 2022; Wei et al. 2022; Bao et al. 2022). To train a powerful model using MIM, one can mask out a part of the foreground of inputs and reconstruct it using the model. Masking out the foreground instead of the background reduces discriminative information in visible information to the model. Reconstructing foreground parts is harder than background and helps the model learn more discriminative information (Wang et al. 2023). Inspired by this phenomenon, we aim to build a more difficult task by adding adversarial perturbations to natural inputs. The adversarial perturbations increase the distance between the input and reconstruction target. If we can reconstruct natural inputs from adversarial examples, it means the features learned by the model are robust against adversarial attacks. In other words, we want the encoder to learn a latent representation that does not carry information concerning adversarial perturbations while enabling the decoder to reconstruct the natural image. Note that simply masking out all foreground does not make sense, as it would render the reconstruction of meaningful content infeasible. From the perspective of IB, there is a bottleneck between the encoder and decoder. Our adversarial pre-training task contains two information sources: the natural data $x$ and the adversarial perturbations $\delta$. As the information flows through the bottleneck, adversarial information from $\delta$ is eliminated. The natural information from $x$ is maintained because of the constraint when reconstructing the target $x$.
Algorithm 1 MIMIR Pre-training
@@
3.3 Design Details
Autoencoder. MIMIR follows the general design of MAE (He et al. 2022), consisting of an encoder $f_{e}$ and a decoder $f_{d}$. As with other autoencoders, the encoder extracts discriminative features $z$ from inputs $x$. The decoder reconstructs original inputs according to the discriminative features. Following ViT (Dosovitskiy et al. 2021), the input $x$ is separated into non-overlapping image patches. We randomly mask out a part of the patches and use the remaining patches as inputs for the following process in the encoder. This random masking process uses uniform distribution to prevent potential sampling bias, such as all foreground information being masked, as it becomes infeasible to find the reconstruction target. Thus, we aim to keep a part of the foreground as a hint for reconstruction. The information of masked content is recorded as mask tokens $m$, not used by the encoder but reserved for later use by the decoder. Each token is a learned vector indicating the presence of a masked patch to be predicted. The mask token is shared by all inputs of the decoder. Like unmasked patches, mask tokens are also assigned corresponding positional embeddings to be in the correct location in the reconstructed image. We emphasize that mask tokens are not used in the encoder part.
To train a ViT, we use the same transformer blocks as ViT to build the encoder. The encoder only processes the visible patches, making training much more efficient. When converting to other architectures, such as ConViT (D’Ascoli et al. 2021), we use corresponding transformer blocks to build the encoder. The decoder accepts the encoded visible image patches and mask tokens as inputs. The decoder is built using the same transformer blocks as the encoder instead of using ViT (Dosovitskiy et al. 2021) transformer blocks for all. Then, the decoder is followed by a fully connected layer, which outputs the same number of patches as the original image.
Adversarial Pre-training Target. The training target is to extract discriminative features from visible image patches by the encoder and then reconstruct the invisible patches by the decoder. Therefore, we need a differentiable measurement to quantify the distance between the original image and the reconstructed results. Following the original MAE (He et al. 2022), this distance is measured by Mean Squared Error (MSE). To create a more difficult reconstruction task, we apply adversarial perturbations $\delta$ on the inputs of the encoder. Thus, the adversarial perturbations are also masked along with the image upon input into the encoder. The decoder reconstructs the original natural inputs by using the latent features $z$ extracted from adversarial examples. The outputs of decoder $x_{re}$ and $x$ are used to calculate the MSE loss, which is further used to optimize the model. Note that the reconstruction differs from the original MAE; we do not use the encoder inputs as reconstruction targets. Formally, the pre-training process (described in Algorithm 1) can be written as follows:
$z=f_{e}(x+\delta),~{}x_{re}=f_{d}(z),$ (4)
$loss_{\text{mse}}=\mathcal{L}_{\text{mse}}(x,x_{re}).$
MI as Penalty. Inspired by IB, we show in Section 3.4 that MI between latent representation and adversarial examples decreases as the accuracy on adversarial examples, i.e., $I(x+\delta,z)$ is decreasing while training. Motivated by this finding, we directly use $I(x+\delta,z)$ as a penalty in our final loss function:
where $\lambda$ is a regularizer for the MI penalty. We use $I(x+\delta,z)$ instead of $I(x,z)$ as a penalty. This is because $x\to x+\delta\to z$ follows the Markov chain since $z$ is extracted from $x+\delta$. According to Data Processing Inequality (DPI) (Beaudry and Renner 2011), $I(x,z)\leq I(x+\delta,z)$. $I(x+\delta,z)$ is closer to $z$ on the Markov chain.
Generating Adversarial Examples. To conduct the adversarial pre-training, we need an attack that finds proper adversarial perturbations $\delta$. As the autoencoder does not provide classification outputs, it is not possible to directly use existing adversarial attacks, such as PGD (Madry et al. 2018). Nevertheless, it is feasible to design a new algorithm to find $\delta$ by maximizing $loss_{\text{mse}}$ in Eq. (4). As the feature $z$ is extracted from only visible image patches, we only attack the visible patches. We do not add any perturbations to mask tokens since the outputs of the autoencoder are only impacted by visible patches. Then, the adversarial pre-training learning objective can be written as: